A survey of the number of children in families in a small town gave the following results.

2 3 1 3 1 2 0 0 1 2 1 0 1 3 1

Does this data provide sufficient evidence that the average number of children per family is less

than 2, at the 10% significance level? Clearly show how you draw a conclusion when you test this

hypothesis. Show calculations of all statistics used.

[20]


1
Expert's answer
2021-10-18T06:35:59-0400

"H_0: \mu = 2 \\

H_1: \mu < 2 \\

n=15 \\

\bar{x} = \frac{2+3+...+3+1}{15} = 1.4 \\

s = \sqrt{\frac{(2-1.4)^2+(3-1.4)^2 +...+(3-1.4)^2+(1-1.4)^2}{15-1}}=1.06"

Test-statistic

"t= \frac{\bar{x}-\mu}{s / \sqrt{n}} \\

t = \frac{1.4-2}{1.06 / \sqrt{15}} = -2.19 \\

α=0.1 \\

df = n-1 = 14"

Critical value

"t_{14,0.10}= -1.35"

Reject H0 if "|t|> |t_{n-1,\u03b1}|"

2.19>1.35

Reject the null hypothesis.

There is sufficient evidence to support that the average number of children per family is less than 2 at a 0.1 level of significance.